# Spatial Scale of Agglomeration and Dispersion Number, Spacing, and the Spatial Extent of Cities

### v7

@@ Appendix E Local advantages
This appendix considers small asymmetries in local regional characteristics, while retaining the circular geographical structure. In particular, we show that if interregional access improves, the regions with higher exogenous local advantage experience population increase under a global dispersion force, and converse is true under a local dispersion force.
@@ E.1   Evaluating the impacts of local characteristics
Consider a spatial model with the indirect utility function $\bm{v}$. Let $a_{i}>0$ denote the innate characteristics of region $i$, and write $\bm{a}=(a_{i})_{i\in\mathcal{I}}$. We assume there is no other source of regional heterogeneity. For example, $a_{i}$ may represent the level of exogenous amenities or productivity in region $i$ [^21] All endogenous mechanisms related to agents’ spatial distribution $\bm{x}$, including endogenous amenities (e.g., Diamond-AER2016), are embedded in the indirect utility function $\bm{v}$ in our framework..
[…]
Example 2.
In Allen-Arkolakis-QJE2014, the indirect utility takes the form
[…]
where $\bm{w}=(w_{i}){i\in\mathcal{I}}$ is the wage vector determined by the market-clearing conditions:
$w_{i}x_{i}=\sum_{j\in\mathcal{I}}\frac{w_{i}^{1-\sigma}\big{(}\overbrace{b_{i}x_{i}^{\alpha}}^{\mathclap{\text{Local productivity spillover (agglomeration force)}}}\big{)}^{\sigma-1}\phi_{ij}}{\sum_{k\in\mathcal{I}}w_{k}^{1-\sigma}\left(b_{k}x_{k}^{\alpha}\right)^{\sigma-1}\phi_{kj}}w_{j}x_{j}\quad\forall i\in\mathcal{I},$ (E.2)
with parameters $\alpha>0$, $\beta<0$, and $\sigma>1$. Here, $u_{i}$ and $b_{i}$ are region-specific primitives interpreted as innate amenities and productivity, respectively. When $\alpha+\beta<0$, the equilibrium is unique. In that case, given an observed population distribution $\bm{x}$, the values of $u_{i}$ and $b_{i}$ that rationalize $\bm{x}$ as an equilibrium are uniquely identified from the equilibrium conditions and interpreted as the regions’ innate characteristics. ∎
When equilibrium is unique, counterfactual analysis proceeds by examining how the population distribution $\bm{x}$ responds to changes in interregional transport costs, holding the calibrated regional characteristics fixed. This assumption that $\bm{a}$ remains unchanged under the counterfactual shocks is central to quantitative predictions. Hence, understanding how $\bm{a}$ influences model outcomes is critical.
@@
$\rho\equiv\sum_{i\in\mathcal{I}}\overbrace{(a_{i}-\bar{a})}^{\mathclap{\text{Regional (dis)advantage}}}\underbrace{(x_{i}(\bm{a})-\bar{x})}_{\mathclap{\text{Population deviation}}}.$ (E.3)
@@ E.2   Formal characterizations
[…]
@@
Assume $N=2$. For the perturbed equilibrium under $\bm{a}=(\bar{a}+\epsilon,\bar{a}-\epsilon)$ with small $\epsilon$,
$\rho=-c\cdot\frac{\omega^{\natural}}{\omega}\quad\text{where}\quad c\equiv 2\epsilon^{2}\frac{\bar{x}}{\bar{a}}.$ (E.6)
@@
Lemma 6 implies the following result on the response of $\rho$ to the increase in $\phi$.
@@
Assume $N=2$. If local regional characteristic is multiplicatively separable as $v_{i}(\bm{x},\bm{a})=a_{i}v_{i}(\bm{x})$ with $v_{i}(\bm{x})$ satisfying Assumption S (i.e., no other exogenous asymmetries),
@@
Under the hypothesis of the claim, $\omega^{\natural}=\frac{\bar{a}}{\bar{v}}(v_{1}(\bar{\bm{x}})-0)=\bar{a}>0$. Then, Eq. E.6 reduces to $\rho^{\prime}(\phi)=\frac{c\bar{a}}{\omega^{2}}\omega^{\prime}(\phi)$ where denotes the differentiation by $\phi$, implying Eq. E.7. ∎
@@ B The Krugman model (Type G)
While the sign of Eq. E.6 is model-dependent, there is a broad connection between the spatial scale of dispersion forces and the sign of $\rho^{\prime}(\phi)$. In Type G models, agglomeration occurs when $\phi$ increases. Reflecting this, for Type G models in the literature, we have $\omega^{\prime}(\phi)>0$ if $\bar{\bm{x}}$ is stable. Converse is true for Type L models, for which dispersion occurs when transport access improves and $\omega^{\prime}(\phi)<0$ if $\bar{\bm{x}}$ is stable. Fig. E.1 shows the curves of $\omega$ for the Krugman model (Type G) and the Redding–Sturm model (Type L), which demonstrates that Types L and G can come to the opposite conclusions in the two-region economy. If a model has only a global dispersion force, its mechanisms strengthen the effects of exogenous local advantages in innate amenities, and the converse is true for a model with only a global dispersion force.
@@
The circular geography. Nonetheless, we have a characterization of the response of $\rho$ for general local characteristics. Now suppose the symmetric circle (Assumption C). Several notations and assumptions are in order. As we have seen in Section 3, the utility elasticity matrix $\mathbf{V}=\frac{\bar{x}}{\bar{v}}[\frac{\partial v_{i}}{x_{j}}(\bar{\bm{x}})]$ at $\bar{\bm{x}}$ is simply represented by the row-normalized proximity matrix $\mathbf{D}$. Suppose $\mathbf{V}=\Omega(\mathbf{D})$ where $\Omega$ is a scalar-valued rational function that is continuous over $[0,1]$ and the interpretation of $\Omega(\mathbf{D})$ is the same as in Section A.2. In an analogous manner, let $\mathbf{A}=\frac{\bar{a}}{\bar{v}}[\frac{\partial v_{i}}{\partial a_{j}}(\bar{\bm{x}})]$ be the utility elasticity matrix with respect to the local characteristic vector $\bm{a}$ of interest, and suppose that $\mathbf{A}$ can be represented by $\mathbf{D}$. Specifically, let $\mathbf{A}=\Omega^{\natural}(\mathbf{D})$ where $\Omega^{\natural}(\cdot)$ is another rational function that is continuous over $[0,1]$. In fact, regional heterogeneities in $\bm{a}$ can be seen as deviations from $\bar{\bm{a}}$, so the effects of such deviations on the utility vector can be evaluated exactly the same manner as population deviations from $\bar{\bm{x}}$ considered in the two-region case. Let
@@
which corresponds to $-\frac{\omega^{\natural}}{\omega}$ in Eq. E.6 for the two-region case.
@@
Suppose Assumption C and assume that $\bar{\bm{x}}$ is stable. Then,
@@
See Section E.4. ∎
The multiplicatively separable case is the special case in which $\Omega^{\natural}(\cdot)=\bar{a}$. For the general case with non-constant $\Omega^{\natural}$, our three model classes are not precisely mapped to Proposition 5 (a) or (b). Still, there is a broad tendency that initial advantages are amplified in Type G models, whereas they are diminished in Type L models. For example, the regional-scale model considered in Redding-Rossi-Hansberg-ARE2017 is Type L. If we consider local productivity parameters as regional characteristic vector $\bm{a}$, then the model satisfies $\delta^{\prime}(\Theta)>0$ for all $\Theta\in(0,1)$ when equilibrium is unique (see Remark 11 in Appendix F). Likewise, the Krugman model is Type G and we show $\delta^{\prime}(\Theta)<0$ for all $\Theta\in(0,1)$ if we consider immobile demand $l_{i}$ as regional characteristics (see Remark 7 in Appendix F).
@@ E.3   Numerical examples
This section provides numerical examples for Proposition 5. To introduce exogenous asymmetry, we multiply the utility in region $1$ by $a_{1}\geq 1$, whereas we let $a_{i}=1$ for all $i\neq 1$. We consider the Krugman model and Allen–Arkolakis model, and basic model parameters are set to be the same as Fig. C.3 and Fig. C.3 except that region $1$ has an exogenous advantage.
Figure 2A reports equilibrium paths of $x_{1}$ for the Allen–Arkolakis model (Type L) under the uniqueness of the equilibrium. The curves depict region $1$’s population share, $x_{1}$, at stable equilibria against $\phi$. Four incremental settings $a_{1}\in\{1.000,1.001,1.005,1.010\}$ are considered, including the baseline case with no location-fixed advantage ($a_{1}=1.000$). We have $\delta^{\prime}(\Theta)>0$ for all $\Theta\in(0,1)$ and see that $x_{1}-\bar{x}>0$ when $a_{1}>1$ and $x_{1}-\bar{x}$ increases as $a_{1}$ increases, which are intuitive. Additionally, $x_{1}-\bar{x}$ decreases as $\phi$ increases. We confirm that $\rho(\phi)>0$ and $\rho^{\prime}(\phi)<0$ for all $\phi$.
Figures 2C and 2B consider the Krugman and Allen–Arkolakis models under a multiplicity of equilibria, respectively. Unlike the Allen–Arkolakis model, the Krugman model admits multiple equilibria for some $\phi$ for any pair of the structural parameters $(\mu,\sigma)$. Proposition 5 correctly predicts the sign of $\rho^{\prime}(\phi)$ for the range of $\phi$ such that $\bar{\bm{x}}$ is stable when $a_{1}=1$; we have $\rho^{\prime}(\phi)>0$ when $\phi\in(0,\phi^{*})$ for the Krugman model, whereas $\rho^{\prime}(\phi)<0$ when $\phi\in(\phi^{**},1)$ for the Allen–Arkolakis model.
@@ C Type G
In Fig. 2C, the definition of $\rho$ is modified for spatial patterns with unpopulated regions. For the range $\phi\in(\phi^{*},\phi^{**})$, $\rho$ is evaluated with respect to the four-centric pattern $(2\bar{x},0,2\bar{x},0,2\bar{x},0,2\bar{x},0)$: $\rho\equiv\sum_{i\in\mathcal{I}(\bm{x})}(x_{i}-2\bar{x})(a_{i}-\bar{a}(\bm{x}))$, where $\mathcal{I}(\bm{x})=\{i\in\mathcal{I}\mid x_{i}>0\}$ is the set of populated regions and $\bar{a}(\bm{x})\equiv\frac{1}{|\mathcal{I}(\bm{x})|}\sum_{i\in\mathcal{I}(\bm{x})}a_{i}$. We define $\rho$ for two-centric pattern $(4\bar{x},0,0,0,4\bar{x},0,0,0)$ similarly. For the transitional phase after ϕsuperscriptitalic-$\phi^{**}$ we let
@@
For Fig. 2B, we employ Eq. E.9 as the definition of $\rho$ for the case $\phi\in(0,\phi^{**})$, i.e., we consider the deviation from the baseline equilibrium ($a_{1}=1$). We observe that $\rho^{\prime}(\phi)<0$ does not necessarily hold true for $\phi\in(0,\phi^{**})$. For instance, $\rho^{\prime}(\phi)>0$ when $\phi$ is small. Nonetheless, for the range of $\phi$ under which $\bar{\bm{x}}$ is stable, $\rho$ decreases in $\phi$, consistent with Proposition 5.
E.4   Proof of Proposition 5
We first derive the analytical expression of $\rho$. The equilibrium condition when all regions are populated is given by
@@ E.4   Proof of Proposition 5
where we make the dependence of $\bm{v}$ on $\bm{a}$ explicit, $\bar{v}(\bm{x},\bm{a})\equiv\sum_{i\in\mathcal{I}}v_{i}(\bm{x},\bm{a})x_{i}$ is the average utility, and $\bm{1}$ is $N$-dimensional all-one vector. The pair $(\bar{\bm{x}},\bar{\bm{a}})$ is a solution to Eq. E.10. Suppose that there is a spatial equilibrium nearby $\bar{\bm{x}}$ when $\bm{a}$ is marginally different from $\bar{\bm{a}}$. Let $\bm{x}(\bm{a})$ denote the perturbed version of $\bar{\bm{x}}$, which is a function in $\bm{a}$. We assume $\bar{\bm{x}}$ is stable so that studying a perturbed version of it makes sense.
@@
where $\mathbf{C}\equiv\mathbf{I}-\frac{1}{N}\bm{1}\bm{1}^{\top}$ is the centering matrix. Let $\mathbf{X}\equiv[\frac{\partial x_{i}}{\partial a_{j}}(\bar{\bm{a}})]$ be the Jacobian matrix of $\bm{x}$ with respect to $\bm{a}$ at $(\bar{\bm{x}},\bar{\bm{a}})$. Then, $\bm{x}(\bm{a})\approx\bar{\bm{x}}+\mathbf{X}(\bm{a}-\bar{\bm{a}})=\bar{\bm{x}}+\mathbf{X}\mathbf{C}\bm{a}$ and thus $\rho=\bm{a}^{\top}\mathbf{C}\mathbf{X}\mathbf{C}\bm{a}$ as $\mathbf{C}\bar{\bm{x}}=\bm{0}$. The implicit function theorem regarding Eq. E.10 at $(\bar{\bm{x}},\bar{\bm{a}})$ gives:
@@
Suppose Assumption C. Let $\mathbf{A}\equiv\frac{\bar{a}}{\bar{v}}[\frac{\partial v_{i}}{\partial a_{j}}]$ be the elasticity matrix of the utility with respect to the local characteristic $\bm{a}$ under consideration, evaluated at $(\bar{\bm{x}},\bar{\bm{a}})$. There is a rational function $\Omega^{\natural}$ that is continuous over $[0,1]$, positive whenever $\bar{\bm{x}}$ is stable, and satisfies $\mathbf{A}=\Omega^{\natural}(\mathbf{D})$. ∎
@@
Suppose $v_{i}(\bm{x},\bm{a})=a_{i}v_{i}(\bm{x})$, where $a_{i}>0$ is the exogenous level of local amenities and $\bm{v}(\bm{x})=(v_{i}(\bm{x}))_{i\in\mathcal{I}}$ is the symmetric component of the utility function (i.e., $\bm{v}(\bm{x})$ satisfies Assumption S). Then, $\mathbf{A}=\bar{a}\mathbf{I}$ and $\Omega^{\natural}(\cdot)=\bar{a}>0$. ∎
Under Assumptions C and A, $\mathbf{X}$ is real, symmetric, and circulant. Thus, the set of eigenvectors of $\mathbf{C}\mathbf{X}\mathbf{C}$ can be chosen as in Lemma 5 (a) because it is a circulant matrix of the same size as $\mathbf{D}$. Let $\{\lambda_{k}\}_{k=0}^{M}$ be the distinct eigenvalues of $\mathbf{C}\mathbf{X}\mathbf{C}$. As $\mathbf{C}\mathbf{X}\mathbf{C}$ is symmetric, it admits the eigenvalue decomposition
@@
where $\tilde{\bm{a}}\equiv(\tilde{a}_{k})$ is the representation of $\bm{a}$ in the new coordinate system $\{\bm{z}_{k}\}$. We can drop $k=0$ because $\lambda_{0}=0$, reflecting that $\bm{z}_{0}=\bm{1}$ represents a uniform increase in $\bm{a}$ and thus does not affect spatial equilibria. All the matrices in Eq. E.15 are circulant and hence shares the same set of eigenvectors. Thus, $\lambda_{k}$ is obtained from Eq. E.15 as follows:
@@
From Eq. E.17, $\rho>0$ for all $\bm{a}$ if all $\{\lambda_{k}\}$ are positive except for $\lambda_{0}=0$. The denominator of Eq. E.19, $\Omega(\Theta_{k})$, must be negative for all $k$ because $\bar{\bm{x}}$ is stable by assumption. Thus, we see that $\rho>0$ if $\Omega^{\natural}(\Theta)>0$ for all $\Theta$ since $\Theta_{k}\in(0,1)$ for all $k\in\mathcal{K}$.
Proposition 5 follows by noting
@@
From Lemma 5, $\{\Theta_{k}\}_{k\in\mathcal{K}}$ are strictly decreasing in $\phi$. Thus, for $\rho^{\prime}(\phi)>0$ ($\rho^{\prime}(\phi)<0$), it is sufficient that $\delta^{\prime}(\Theta)<0$ ($\delta^{\prime}(\Theta)>0$) for all $\Theta$ such that $\Omega(\Theta)<0$.

### v8

@@ C Type LG
Note: There are no asymmetries in regional characteristics except for geographic accessibility. The transport cost between every consecutive pair of regions is uniform. The level of transport cost monotonically decreases from top to bottom. Panels (A) and (B) consider a line segment with 65 locations. Panel (B) adapted from Ikeda-et-al-IJET2017 considers 17 locations; see the original paper for an extensive discussion.
@@ Appendix E Local advantages
This appendix introduces small region-specific asymmetries within a circular geography. We show that improved interregional access raises population in regions with greater exogenous advantage under global dispersion forces, but lowers it under local dispersion forces.
@@ E.1  Evaluating the impacts of local characteristics
Consider a spatial model with the indirect utility function $\bm{v}$. Let $a_{i}>0$ denote the innate characteristics of region $i$, and write $\bm{a}=(a_{i})_{i\in\mathcal{I}}$. For example, $a_{i}$ may represent the level of exogenous amenities or productivity in region ii [^25] All endogenous mechanisms related to agents’ spatial distribution $\bm{x}$, including endogenous amenities (e.g., Diamond-AER2016), are embedded in the indirect utility function $\bm{v}$ in our framework.. All regions are perfectly symmetric if $a_{i}=\bar{a}>0$ for all $i$. In this case, the uniform population distribution $\bar{\bm{x}}$ is an equilibrium.
When equilibrium is unique, counterfactual analysis proceeds by examining how the population distribution $\bm{x}$ responds to changes in interregional transport costs, holding the calibrated regional characteristics fixed. This assumption that $\bm{a}$ remains unchanged under the counterfactual shocks is central to quantitative predictions. Hence, it is important to understand how $\bm{a}$ influences model outcomes.
@@
[…]
@@ E.2  Formal characterizations
Two regions. Assume that two regions have the same local characteristics and assume that $\bar{\bm{x}}=(\bar{x},\bar{x})$ is stable. Consider a marginal regional asymmetry of the form $\bm{a}=(\bar{a}+\epsilon,\bar{a}-\epsilon)$ with small $\epsilon$, so that $\bar{\bm{x}}$ is perturbed to a new equilibrium $\bm{x}=(\bar{x}+\xi,\bar{x}-\xi)$ with small $\xi$. We can assume that $\xi>0$ and $\epsilon>0$. Then, by definition, we have
@@
Assume $N=2$. For the perturbed equilibrium under $\bm{a}=(\bar{a}+\epsilon,\bar{a}-\epsilon)$ with small $\epsilon$, we have $\rho=-c\cdot\frac{\omega^{\natural}}{\omega}$ where $c\equiv 2\epsilon^{2}\frac{\bar{x}}{\bar{a}}$. ∎
@@
Lemma˜6 implies the following result on the response of $\rho$ to the increase in $\phi$.
@@
Assume $N=2$. If local regional characteristic is multiplicatively separable as $v_{i}(\bm{x},\bm{a})=a_{i}v_{i}(\bm{x})$ with $v_{i}(\bm{x})$ satisfying Assumption˜S (i.e., no other exogenous asymmetries),
@@
Under the hypothesis of the claim, $\omega^{\natural}=\frac{\bar{a}}{\bar{v}}(v_{1}(\bar{\bm{x}})-0)=\bar{a}>0$. Then, $\rho=-c\cdot\frac{\omega^{\natural}}{\omega}=-\frac{c\bar{a}}{\omega}$ and $\rho^{\prime}(\phi)=\frac{c\bar{a}}{\omega^{2}}\omega^{\prime}(\phi)$ where denotes the differentiation by $\phi$, implying Eq.˜E.4. ∎
@@ B The Krugman model (Type G)
There is a broad connection between the spatial scale of dispersion forces and the sign of $\rho^{\prime}(\phi)$. In Type G models, agglomeration occurs when $\phi$ increases. Reflecting this, for Type G models in the literature, we have $\omega^{\prime}(\phi)>0$ if $\bar{\bm{x}}$ is stable. Converse is true for Type L models, for which dispersion occurs when transport access improves and $\omega^{\prime}(\phi)<0$ if $\bar{\bm{x}}$ is stable. Fig.˜E.1 shows the curves of $\omega$ for the Krugman model (Type G) and the Redding–Sturm model (Type L), which demonstrates that Types L and G can come to the opposite conclusions in the two-region economy. If a model has only a global dispersion force, its mechanisms strengthen the effects of exogenous local advantages in innate amenities, and the converse is true for a model with only a global dispersion force.
@@
The circular geography. Nonetheless, we have a characterization of the response of $\rho$ for general local characteristics. Now suppose the symmetric circle (Assumption˜C). Several notations and assumptions are in order. As we have seen in Section˜3, the utility elasticity matrix $\mathbf{V}=\frac{\bar{x}}{\bar{v}}[\frac{\partial v_{i}}{x_{j}}(\bar{\bm{x}})]$ at $\bar{\bm{x}}$ is simply represented by the row-normalized proximity matrix $\mathbf{D}$. Suppose $\mathbf{V}=\Omega(\mathbf{D})$ where $\Omega$ is a scalar-valued rational function that is continuous over $[0,1]$ and the interpretation of $\Omega(\mathbf{D})$ is the same as in Section˜A.3. In an analogous manner, let $\mathbf{A}=\frac{\bar{a}}{\bar{v}}[\frac{\partial v_{i}}{\partial a_{j}}(\bar{\bm{x}})]$ be the utility elasticity matrix with respect to the local characteristic vector $\bm{a}$ of interest, and suppose that $\mathbf{A}$ can be represented by $\mathbf{D}$. Specifically, let $\mathbf{A}=\Omega^{\natural}(\mathbf{D})$ where $\Omega^{\natural}(\cdot)$ is another rational function that is continuous over $[0,1]$. In fact, regional heterogeneities in $\bm{a}$ can be seen as deviations from $\bar{\bm{a}}$, so the effects of such deviations on the utility vector can be evaluated exactly the same manner as population deviations from $\bar{\bm{x}}$ considered in the two-region case. Let
@@
which corresponds to $\frac{\omega^{\natural}}{\omega}$ for the two-region case.
@@
Suppose Assumption˜C and assume that $\bar{\bm{x}}$ is stable. Then,
@@
See Section˜E.4. ∎
The multiplicatively separable case is the special case in which $\Omega^{\natural}(\cdot)=\bar{a}$. For the general case with non-constant $\Omega^{\natural}$, our three model classes are not precisely mapped to Proposition˜5 (a) or (b). Still, there is a broad tendency that initial advantages are amplified in Type G models, whereas they are diminished in Type L models. For example, the regional-scale model considered in Redding-Rossi-Hansberg-ARE2017 is Type L. If we consider local productivity parameters as regional characteristic vector $\bm{a}$, then the model satisfies $\delta^{\prime}(\Theta)>0$ for all $\Theta\in(0,1)$ when equilibrium is unique (see Remark˜9 in Appendix˜F). Likewise, the Krugman model is Type G and we show $\delta^{\prime}(\Theta)<0$ for all $\Theta\in(0,1)$ if we consider immobile demand $l_{i}$ as regional characteristics (see Remark˜6 in Appendix˜F).
@@ E.3  Numerical examples
This section provides numerical examples for Proposition˜5. To introduce exogenous asymmetry, we multiply the utility in region $1$ by $a_{1}\geq 1$, whereas we let $a_{i}=1$ for all $i\neq 1$. We consider the Krugman model and Allen–Arkolakis model, and basic model parameters are set to be the same as Fig.˜C.3 and Fig.˜C.3 except that region $1$ has an exogenous advantage.
Figure˜2A reports equilibrium paths of $x_{1}$ for the Allen–Arkolakis model (Type L) under the uniqueness of the equilibrium. The curves depict region $1$’s population share, $x_{1}$, at stable equilibria against $\phi$. Four incremental settings $a_{1}\in\{1.000,1.001,1.005,1.010\}$ are considered, including the baseline case with no location-fixed advantage ($a_{1}=1.000$). We have $\delta^{\prime}(\Theta)>0$ for all $\Theta\in(0,1)$ and see that $x_{1}-\bar{x}>0$ when $a_{1}>1$ and $x_{1}-\bar{x}$ increases as $a_{1}$ increases, which are intuitive. Additionally, $x_{1}-\bar{x}$ decreases as $\phi$ increases. We confirm that $\rho(\phi)>0$ and $\rho^{\prime}(\phi)<0$ for all $\phi$.
Figures˜2C and 2B consider the Krugman and Allen–Arkolakis models under a multiplicity of equilibria, respectively. Unlike the Allen–Arkolakis model, the Krugman model admits multiple equilibria for some $\phi$ for any pair of the structural parameters $(\mu,\sigma)$. Proposition˜5 correctly predicts the sign of $\rho^{\prime}(\phi)$ for the range of $\phi$ such that $\bar{\bm{x}}$ is stable when $a_{1}=1$; we have $\rho^{\prime}(\phi)>0$ when $\phi\in(0,\phi^{*})$ for the Krugman model, whereas $\rho^{\prime}(\phi)<0$ when $\phi\in(\phi^{**},1)$ for the Allen–Arkolakis model.
@@ C Type G
In Fig.˜2C, the definition of $\rho$ is modified for spatial patterns with unpopulated regions. For the range $\phi\in(\phi^{*},\phi^{**})$, $\rho$ is evaluated with respect to the four-centric pattern $(2\bar{x},0,2\bar{x},0,2\bar{x},0,2\bar{x},0)$: $\rho\equiv\sum_{i\in\mathcal{I}(\bm{x})}(x_{i}-2\bar{x})(a_{i}-\bar{a}(\bm{x}))$, where $\mathcal{I}(\bm{x})=\{i\in\mathcal{I}\mid x_{i}>0\}$ is the set of populated regions and $\bar{a}(\bm{x})\equiv\frac{1}{|\mathcal{I}(\bm{x})|}\sum_{i\in\mathcal{I}(\bm{x})}a_{i}$. We define $\rho$ for two-centric pattern $(4\bar{x},0,0,0,4\bar{x},0,0,0)$ similarly. For the transitional phase after $\phi^{**}$ we let
@@
For Fig.˜2B, we employ Eq.˜E.6 as the definition of $\rho$ for the case $\phi\in(0,\phi^{**})$, i.e., we consider the deviation from the baseline equilibrium ($a_{1}=1$). We observe that $\rho^{\prime}(\phi)<0$ does not necessarily hold true for $\phi\in(0,\phi^{**})$. For instance, $\rho^{\prime}(\phi)>0$ when $\phi$ is small. Nonetheless, for the range of $\phi$ under which $\bar{\bm{x}}$ is stable, $\rho$ decreases in $\phi$, consistent with Proposition˜5.
E.4  Proof of Proposition˜5
We derive the analytical expression of $\rho$. If all regions are populated in equilibrium, we have
@@ E.4  Proof of Proposition˜5
where we make the dependence of $\bm{v}$ on $\bm{a}$ explicit, $\bar{v}(\bm{x},\bm{a})\equiv\sum_{i\in\mathcal{I}}v_{i}(\bm{x},\bm{a})x_{i}$ is the average utility, and $\bm{1}$ is $N$-dimensional all-one vector. The pair $(\bar{\bm{x}},\bar{\bm{a}})$ is a solution to Eq.˜E.7. Suppose that there is a spatial equilibrium nearby $\bar{\bm{x}}$ when $\bm{a}$ is marginally different from $\bar{\bm{a}}$. Let $\bm{x}(\bm{a})$ denote the perturbed version of $\bar{\bm{x}}$, which is a function in $\bm{a}$. We assume $\bar{\bm{x}}$ is stable so that studying a perturbed version of it makes sense.
@@
where $\mathbf{C}\equiv\mathbf{I}-\frac{1}{N}\bm{1}\bm{1}^{\top}$ is the centering matrix. Let $\mathbf{X}\equiv[\frac{\partial x_{i}}{\partial a_{j}}(\bar{\bm{a}})]$ be the Jacobian matrix of $\bm{x}$ with respect to $\bm{a}$ at $(\bar{\bm{x}},\bar{\bm{a}})$. Then, $\bm{x}(\bm{a})\approx\bar{\bm{x}}+\mathbf{X}(\bm{a}-\bar{\bm{a}})=\bar{\bm{x}}+\mathbf{X}\mathbf{C}\bm{a}$ and thus $\rho=\bm{a}^{\top}\mathbf{C}\mathbf{X}\mathbf{C}\bm{a}$ as $\mathbf{C}\bar{\bm{x}}=\bm{0}$. The implicit function theorem regarding Eq.˜E.7 at $(\bar{\bm{x}},\bar{\bm{a}})$ gives:
@@
Suppose Assumption˜C. Let $\mathbf{A}\equiv\frac{\bar{a}}{\bar{v}}[\frac{\partial v_{i}}{\partial a_{j}}]$ be the elasticity matrix of the utility with respect to the local characteristic $\bm{a}$, evaluated at $(\bar{\bm{x}},\bar{\bm{a}})$. There is a rational function $\Omega^{\natural}$ that is continuous over $[0,1]$, positive whenever $\bar{\bm{x}}$ is stable, and satisfies $\mathbf{A}=\Omega^{\natural}(\mathbf{D})$. ∎
@@
Suppose $v_{i}(\bm{x},\bm{a})=a_{i}v_{i}(\bm{x})$, where $a_{i}>0$ is the exogenous level of local amenities and $\bm{v}(\bm{x})=(v_{i}(\bm{x}))_{i\in\mathcal{I}}$ is the symmetric component of the utility function (i.e., $\bm{v}(\bm{x})$ satisfies Assumption˜S). Then, $\mathbf{A}=\bar{a}\mathbf{I}$ and $\Omega^{\natural}(\cdot)=\bar{a}>0$. ∎
Under Assumptions C and A, $\mathbf{X}$ is real, symmetric, and circulant. Thus, the set of eigenvectors of $\mathbf{C}\mathbf{X}\mathbf{C}$ can be chosen as in Lemma˜5 (a) because it is a circulant matrix of the same size as $\mathbf{D}$. Let $\{\lambda_{k}\}_{k=0}^{M}$ be the distinct eigenvalues of $\mathbf{C}\mathbf{X}\mathbf{C}$. As $\mathbf{C}\mathbf{X}\mathbf{C}$ is symmetric, it admits the eigenvalue decomposition
@@
where $\tilde{\bm{a}}\equiv(\tilde{a}_{k})$ is the representation of $\bm{a}$ in the new coordinate system $\{\bm{z}_{k}\}$. We can drop $k=0$ because $\lambda_{0}=0$, reflecting that $\bm{z}_{0}=\bm{1}$ represents a uniform increase in $\bm{a}$ and thus does not affect spatial equilibria. All the matrices in Eq.˜E.12 are circulant and hence shares the same set of eigenvectors. Thus, $\lambda_{k}$ is obtained from Eq.˜E.12 as follows:
@@
From Eq.˜E.14, $\rho>0$ for all $\bm{a}$ if all $\{\lambda_{k}\}$ are positive except for $\lambda_{0}=0$. The denominator of Eq.˜E.16, $\Omega(\Theta_{k})$, must be negative for all $k$ because $\bar{\bm{x}}$ is stable by assumption. Thus, we see that $\rho>0$ if $\Omega^{\natural}(\Theta)>0$ for all $\Theta$ since $\Theta_{k}\in(0,1)$ for all $k\in\mathcal{K}$.
Proposition˜5 follows by noting
@@
From Lemma˜5, $\{\Theta_{k}\}_{k\in\mathcal{K}}$ are strictly decreasing in $\phi$. Thus, for $\rho^{\prime}(\phi)>0$ ($\rho^{\prime}(\phi)<0$), it is sufficient that $\delta^{\prime}(\Theta)<0$ ($\delta^{\prime}(\Theta)>0$) for all $\Theta$ such that $\Omega(\Theta)<0$.